OpenStudy (anonymous):

How do i solve this?? x+11/x+3 multiplied by x^2-9/x^2+20x+99

OpenStudy (helder_edwin):

it the same as the problem you wrote before you have to factor the second fraction

OpenStudy (anonymous):

how do you factor x^-9 and x^2+20x + 99?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

yeah isn't how you answer :c try factor them~

OpenStudy (helder_edwin):

come on! you have to learn this! look at the previous excercise @nphuongsun93 did 4 u! how do you factor \[ \Large x^2-9=(x\qquad)(x\qquad) \] ?

OpenStudy (anonymous):

would it be x-3 and x +3

OpenStudy (helder_edwin):

yes!

OpenStudy (helder_edwin):

now factor \[ \Large x^2+20x+99=(x+\qquad)(x+\qquad) \]

OpenStudy (anonymous):

okay.give me a sec to work it out:)

OpenStudy (helder_edwin):

remember! you are looking for two numbers such that multiplied =99 and added =20

OpenStudy (anonymous):

okay so i got. (x+9) and (x+11)

OpenStudy (anonymous):

\[x^2+20x+99\] \[x^2+11x+9x+99\] \[x(x+11)+9(x+11)\] \[(x+9)(x+11)\]

OpenStudy (helder_edwin):

good work!

OpenStudy (helder_edwin):

Ok now YOU have \[ \Large \frac{x+11}{x+3}\cdot\frac{x^2-9}{x^2+20x+99}= \] \[ \Large \frac{x+11}{x+3}\cdot\frac{(x-3)(x+3)}{(x+9)(x+11)} \] what do you get?

OpenStudy (anonymous):

x-3 and x+9 ?

OpenStudy (helder_edwin):

and?

OpenStudy (helder_edwin):

you are working on fractions!

OpenStudy (anonymous):

x-3/ x+9 is that how i should write it ?

OpenStudy (helder_edwin):

yes! \[ \Large \frac{x-3}{x+9} \]

OpenStudy (anonymous):

okay!:) thanks for helping!

OpenStudy (helder_edwin):

you are welcome